the sum of the numerator and denominator of fraction is 12. if the denominator is increased by 3, the fractions simplifies to 1/2. what is the original fraction?
Answer by macston(5194)

.
N=numerator; D=denominator
N+D=12
D=12-N
.
2N=D+3
2N=(12-N)+3
3N=15
N=5
.
D=12-N=12-5=7
D/N=5/7
.
ANSWER: The original fraction was 5/7.